. An experimentalist has measured the u-velocity component of a steady, two- imensional flow field. It is approximated by u 3x2y x +10 It is also known that the v-velocity is zero along the line y-0. a) Find an expression for the v-velocity in the entire field b) Find an expression for the streamfunction, v, for this flow c) Determine the location of any stagnation points in the flow (stagnation means -0) d) Calculate the acceleration field (a and ay) for this flow field.
Answer #1

(a): Apply continuity equation for 2D flow as follows: Ou avo * oy 818 818 818 = (3x²y+x+10) =-(6xy+1) =-bry - 1 â=(-6x-1)ây(b): Write the expression for the stream function for incompressible, irrotational, and steady 2D flow as follows: Oy ov OY =(c): At stagnation point, v=0 -3.xy? - y=0 y (3xy +1)=0 y=0 At stagnation point, u=0 3r- x0+x+10=0 x+10=0 r=-10 Thus, the loc